# BranchGRPO: Stable and Efficient GRPO with Structured Branching in Diffusion Models

### v4

@@ Abstract
Recent progress in aligning image and video generative models with Group Relative Policy Optimization (GRPO) has improved human preference alignment, but existing variants remain inefficient due to sequential rollouts and large numbers of sampling steps, unreliable credit assignment: sparse terminal rewards are uniformly propagated across timesteps, failing to capture the varying criticality of decisions during denoising. In this paper, we present BranchGRPO, a method that restructures the rollout process into a branching tree, where shared prefixes amortize computation and pruning removes low-value paths and redundant depths. BranchGRPO introduces three contributions: (1) a branching scheme that amortizes rollout cost through shared prefixes while preserving exploration diversity; (2) a reward fusion and depth-wise advantage estimator that transforms sparse terminal rewards into dense step-level signals; and (3) pruning strategies that cut gradient computation but leave forward rollouts and exploration unaffected. On HPDv2.1 image alignment, BranchGRPO improves alignment scores by up to 16% over DanceGRPO, while reducing per-iteration training time by nearly 55%. A hybrid variant, BranchGRPO-Mix, further accelerates training to 4.7× faster than DanceGRPO without degrading alignment. On WanX video generation, it further achieves higher Video-Align scores with sharper and temporally consistent frames compared to DanceGRPO. Codes are available at BranchGRPO.
@@ 1 Introduction
Within RLHF, Group Relative Policy Optimization (GRPO) has shown strong stability and scalability across text-to-image and text-to-video tasks (Liu et al., 2025a; Xue et al., 2025). However, when applied to diffusion and flow-matching models, current GRPO variants still face two fundamental bottlenecks: (1) Inefficiency. Standard GRPO adopts a sequential rollout design, where each trajectory must be independently sampled under both the old and new policies. This incurs $O(N\cdot T)$ complexity with denoising steps $T$ and group size $N$, leading to significant computational redundancy and limiting scalability in large-scale image and video generation tasks. (2) Sparse rewards. Existing methods assign a single terminal reward uniformly across all steps, neglecting informative signals from intermediate states. This uniform propagation leads to unreliable credit assignment and high-variance gradients, raising the central question: how can we attribute sparse outcome rewards to the specific denoising steps that truly shape final quality?
@@
The tree structure further enables a novel reward fusion with depth-wise advantage estimation. Instead of uniformly propagating a single terminal reward, BranchGRPO aggregates leaf rewards and propagates them backward with depth-wise normalization, producing finer-grained step-level advantages. In addition, width- and depth-pruning strategies remove redundant branches during backpropagation, accelerating training and reallocating computation toward promising regions of the trajectory space.
@@
  • We introduce BranchGRPO, a tree-structured rollout framework that replaces independent sequential sampling with branching trajectories. By inserting split points during denoising, trajectories expand into multiple sub-trajectories while sharing early prefixes, which reduces redundancy and improves exploration efficiency.
@@
We validate BranchGRPO on both text-to-image and image-to-video alignment tasks, demonstrating its effectiveness and generality across modalities.
@@ 2 Related Work
Diffusion models (Ho et al., 2020; Rombach et al., 2022) and flow matching models (Lipman et al., 2022; Liu et al., 2022) have become dominant paradigms for visual generation due to their strong theoretical foundations and ability to generate high-quality content efficiently. While pretraining establishes the generative prior, aligning outputs with nuanced human preferences requires reinforcement learning from human feedback (RLHF). In natural language processing, RLHF has proven highly successful for aligning large language models (LLMs) (Ouyang et al., 2022; Christiano et al., 2017), where methods such as PPO and GRPO enable stable preference-driven post-training. These successes have inspired adaptation of RLHF to vision.
@@ 3.2 Branch Rollout Algorithm
Given a prompt, BranchGRPO reformulates denoising into a tree-structured process. We align terminology with tree search: (i) depth $T$ denotes the number of denoising steps; (ii) width $w$ is the number of completed trajectories (leaves); (iii) branching steps $\mathcal{B}$ indicate split timesteps; (iv) branch correlation $s$ controls the correlation among child noises; (v) branching factor $K$ is the number of children per split.
@@
Unlike prior GRPO variants such as DanceGRPO and FlowGRPO that rely on sequential rollouts, where each trajectory is sampled independently from start to finish, BranchGRPO reorganizes the process into a tree-structured rollout (Figure 1). For each prompt, we initialize a root node with a same initial noise $z_{0}\sim\mathcal{N}(0,I)$ and then denoise step by step along the reverse SDE. At designated split steps $\mathcal{B}$, the current state expands into $K$ children, producing multiple sub-trajectories that share early prefixes but diverge afterward. The branching is achieved by injecting stochastic perturbations into the SDE transition, with a hyperparameter $s$ controlling the correlation strength among child noises. This mechanism balances exploration diversity and stability while keeping the marginal distribution unchanged. The rollout continues until reaching the maximum depth $T$, at which point all leaves are collected for reward evaluation.
Formally, following Xue et al. (2025), the reverse-time dynamics can be written in the unified SDE form:
@@
$z_{i+1}^{(b)}=\mu_{\theta}(z_{i},t_{i})\;+\;\varepsilon_{t_{i}}g_{t_{i}}\sqrt{h_{i}}\,\xi_{b},\qquad\xi_{b}=\frac{\xi_{0}+s\,\eta_{b}}{\sqrt{1+s^{2}}},\quad b=1,\dots,K,$ (2)
@@
Figure 3: Left: branching rollouts generate diverse leaves. Middle: leaf rewards are fused upward by path-probability weighting. Right: depth-wise normalization and pruning yield dense advantages and reduce cost.
@@ 3.3 Reward Fusion and Depth-wise Advantage Estimation
Branch rollouts construct a trajectory tree where all leaves share ancestral prefixes. This shared structure enables internal node values to be represented by their descendant rewards, naturally allowing backward propagation of leaf signals.
In contrast, existing GRPO variants update every step using only a single terminal reward, ignoring intermediate states and resulting in unreliable credit assignment with high variance (Fig. 3). BranchGRPO mitigates this limitation by propagating leaf rewards upward, fusing them through path-probability weighting, and applying depth-wise normalization to produce dense step-level advantages. The optimization objective remains the clipped GRPO loss defined on tree edges:
where $\mathcal{E}$ denotes tree edges and $\rho_{e}(\theta)$ the importance ratio.
@@
Directly averaging leaf rewards may be dominated by rare noisy trajectories. We instead apply a weighted fusion: for internal node $n$ with descendant leaves $\mathcal{L}(n)$,
@@
Here $p_{\text{beh}}$ is the behavior policy and $\beta$ controls concentration. $\beta=0$ reduces to uniform averaging; $\beta=1$ recovers self-normalized importance weighting aligned with the behavior distribution. Uniform averaging is robust to log-prob calibration errors and encourages exploration by retaining low-probability leaves, but introduces variance when branches are many. Path-weighting reduces variance and stabilizes updates, though it may over-concentrate on high-likelihood leaves in deep trees. We empirically compare both variants in Sec. 4.4.
@@ 3.4 Pruning Strategies
Width Pruning.
After computing rewards and advantages for all leaves $\mathcal{L}$, we restrict gradient updates to a subset of them. We investigate two modes. The first, Parent-Top1, retains the child with the higher reward from each parent at the last branching step. This strategy roughly halves gradient computation while ensuring coverage of all parents, yielding stable but slightly less diverse updates. The second, Extreme selection, preserves both the globally best and worst $b$ leaves by reward score. This explicitly maintains strong positive and negative signals, which may enhance exploration but also increase variance.
Depth Pruning.
Branch rollouts generate dense rewards across all denoising steps, but computing gradients at every depth remains costly. To improve efficiency, we introduce depth pruning, which skips gradient computation at selected timesteps . Concretely, we maintain a set of pruned depths $\mathcal{D}$ and ignore gradients from nodes at these layers. To prevent permanently discarding certain steps, we adopt a sliding window schedule: the pruned depths gradually shift toward later timesteps as training progresses, until reaching a predefined maximum depth. Formally, pruning is active throughout training, and at fixed intervals the window slides one step deeper until the stop depth is reached.
@@
17:   Depth-wise normalization: standardize per depth, assign edge advantages
18:   Pruning (optional): select nodes for backprop only
19:   Compute clipped GRPO loss $J(\theta)$
20:  end for
21:  Update policy: gradient ascent on $J(\theta)$
@@ 4.1 Experiment Setup
We evaluate BranchGRPO on HPDv2.1 (Wu et al., 2023) (103k training and 400 balanced test prompts). The backbone is FLUX.1-Dev, with rollouts augmented by branching and process-dense rewards; we also study a pruned variant (BranchGRPO-Pruning). Baselines include DanceGRPO and MixGRPO under identical settings. We report efficiency (NFE, iteration time, GPU-hours) and quality (HPS-v2.1, PickScore (Kirstain et al., 2023), ImageReward (Xu et al., 2023)).
@@ 4.2 Implementation Details
We set the tree depth to $d=20$ and the branch factor to $b=2$, yielding $16$ leaves per rollout before pruning. The branching steps $\mathcal{B}$ use three presets: Dense $(0,3,6,9)$ as the default, Mixed $(0,5,15,20)$, and Sparse $(0,10,13,19)$. The branch correlation sweeps $s\in\{0,1,2,4,8\}$. Training runs for 300 optimizer steps with gradient accumulation $g=12$ and per-GPU batch size =2=2, on $16\times$ NVIDIA H200 GPUs. Optimization uses AdamW (learning rate $1\times 10^{-5}$, weight decay $1\times 10^{-4}$) with bf16 precision and EMA weights stored on CPU. All GRPO-related hyperparameters are kept identical across methods, with full details deferred to the supplementary material.
@@
FLUX - - - 0.313 0.227 1.112
DanceGRPO(tf=1.0) 20 20 698 0.360 0.229 1.189
DanceGRPO(tf=0.6) 20 12 469 0.353 0.228 1.219
MixGRPO (20,5) 20 5 289 0.359 0.228 1.211
BranchGRPO 13.68 13.68 493 0.363 0.229 1.233
BranchGRPO-WidPru 13.68 8.625 314 0.364 0.230 1.300
BranchGRPO-DepPru 13.68 8.625 314 0.369 0.231 1.319
BranchGRPO-Mix 13.68 4.25 148 0.363 0.230 1.290
@@ 4.3 Main Results
Table 1 compares efficiency and alignment quality. BranchGRPO achieves superior alignment while substantially reducing cost: for example, BranchGRPO-DepPruning improves HPS-v2.1 from 0.360 (DanceGRPO) to 0.369 (+2.5%) and attains the best PickScore (0.231) and ImageReward (1.319). Meanwhile, BranchGRPO-Mix accelerates training to 148s per iteration—over $4.7\times$ faster than DanceGRPO(tf=1.0)—while maintaining competitive alignment scores. MixGRPO also reduces computation (289s) but its alignment scores plateau at DanceGRPO levels.
The reward curves in Figure 1 further highlight stability. DanceGRPO(tf=0.6) suffers from high-variance fluctuations, while the full-timestep variant converges more smoothly but at much higher cost. BranchGRPO combines the advantages of both, yielding faster early reward growth and smoother convergence. Qualitative comparisons in Figure 4 show that our generations exhibit sharper details and stronger semantic alignment compared with Flux and DanceGRPO.
@@ 4.4 Ablation Studies
Figure 5: Ablation studies of BranchGRPO. Moderate branch correlation, early and denser splits improve reward growth; path-weighted fusion enhances stability; depth pruning achieves the best final reward; and the hybrid ODE–SDE schedule provides the fastest training speed while remaining stable.
Reward Fusion Strategies. Figure 5(d) compares uniform averaging with path-probability weighting under identical training settings. Uniform averaging shows higher variance and a clear late-stage plateau, whereas path-weighted fusion delivers consistently higher and more stable rewards throughout training. This empirically supports Sec. 3.3: uniform averaging is exploration-friendly but noisy, while path weighting aligns credit assignment with the behavior distribution and improves convergence at no extra cost.
Pruning Strategies. Figure 5(e) evaluates pruning applied after depth-wise normalization and only for backprop. Depth pruning (sliding window over denoising steps) achieves the best final reward, consistently surpassing the no-pruning baseline by reallocating compute to informative early steps while preserving forward sampling diversity. Width pruning (Parent-Top1) halves gradient updates and yields the smoothest, lowest-variance curve, but its final reward is slightly below depth pruning. Width pruning (Extreme-$b$) keeps the globally best and worst $b$ leaves; it injects strong positive/negative signals and remains competitive at the end, but exhibits higher variance than the other settings.

### v5

@@ Abstract
Recent progress in aligning image and video generative models with Group Relative Policy Optimization (GRPO) has improved human preference alignment, but existing variants remain inefficient due to sequential rollouts and large numbers of sampling steps, unreliable credit assignment, as sparse terminal rewards are uniformly propagated across timesteps, failing to capture the varying criticality of decisions during denoising. In this paper, we present BranchGRPO, a method that restructures the rollout process into a branching tree, where shared prefixes amortize computation and pruning removes low-value paths and redundant depths. BranchGRPO introduces three contributions: (1) a branching scheme that amortizes rollout cost through shared prefixes while preserving exploration diversity; (2) a reward fusion and depth-wise advantage estimator that transforms sparse terminal rewards into dense step-level signals; and (3) pruning strategies that cut gradient computation but leave forward rollouts and exploration unaffected. On HPSv2.1 image alignment, BranchGRPO improves alignment scores by up to 16% over DanceGRPO, while reducing per-iteration training time by nearly 55%. A hybrid variant, BranchGRPO-Mix, further accelerates training to 4.7× faster than DanceGRPO without degrading alignment. On WanX video generation, it further achieves higher motion quality reward with sharper and temporally consistent frames.
@@ 1 Introduction
Within RLHF, Group Relative Policy Optimization (GRPO) has shown strong stability and scalability across text-to-image and text-to-video tasks (Liu et al., 2025a; Xue et al., 2025). However, when applied to diffusion and flow-matching models, current GRPO variants still face two fundamental bottlenecks: (1) Inefficiency. Standard GRPO adopts a sequential rollout design, where each trajectory must be independently sampled under both the old and new policies. This incurs $O(N\cdot T)$ complexity with denoising steps $T$ and group size $N$, leading to significant computational redundancy and limiting scalability in large-scale image and video generation tasks. (2) Sparse rewards. Existing methods assign a single terminal reward uniformly across all denoising steps, neglecting informative signals from intermediate states. This uniform propagation leads to unreliable credit assignment and high-variance gradients, raising the central question: how can we attribute sparse outcome rewards to the specific denoising steps that truly shape final quality?
@@
The tree structure further enables a novel reward fusion with depth-wise advantage estimation. Instead of uniformly propagating a single terminal reward, BranchGRPO aggregates leaf rewards and propagates them backward with depth-wise normalization, producing finer-grained step-level advantages. In addition, width- and depth-pruning strategies remove redundant branches and depths during backpropagation, accelerating training and reallocating computation toward promising regions of the trajectory space.
We validate BranchGRPO on both text-to-image and image-to-video alignment tasks, demonstrating its effectiveness and generality across modalities. In addition, we verify the scaling law of BranchGRPO, larger group sizes consistently lead to better alignment performance.
@@
  • We introduce BranchGRPO, a tree-structured GRPO training paradigm. It replaces independent sequential rollouts with branching during denoising, reusing shared prefixes to amortize compute and broaden exploration, thereby improving efficiency and scalability.
@@ 2 Related Work
Diffusion models (Ho et al., 2020; Rombach et al., 2022) and flow matching models (Lipman et al., 2022; Liu et al., 2022) have become dominant paradigms for visual generation due to their strong theoretical foundations and ability to generate high-quality content efficiently. While pretraining establishes the generative prior, aligning outputs with nuanced human preferences requires reinforcement learning from human feedback (RLHF). In natural language processing, RLHF has proven highly successful for aligning large language models (LLMs) (Ouyang et al., 2022; Christiano et al., 2017; Lu et al., 2025), where methods such as PPO and GRPO enable stable preference-driven post-training. These successes have inspired adaptation of RLHF to vision.
@@ 3.2 Branch Rollout Algorithm
Given a prompt, BranchGRPO reformulates denoising into a tree-structured process. We align terminology with tree search: (i) depth $T$ denotes the number of denoising steps; (ii) width $w$ is the number of completed trajectories (leaves); (iii) branching steps $\mathcal{B}$ indicate split timesteps; (iv) branch correlation $s$ controls the diversity among child nodes; (v) branching factor $K$ is the number of children per split.
@@
Unlike prior GRPO variants such as DanceGRPO and FlowGRPO that rely on sequential rollouts, where each trajectory is sampled independently from start to finish, BranchGRPO reorganizes the process into a tree-structured rollout (Figure 1). For each prompt, we initialize a root node with a same initial noise $z_{0}\sim\mathcal{N}(0,I)$ and then denoise step by step along the reverse SDE. At designated split steps $\mathcal{B}$, the current state expands into $K$ children, producing multiple sub-trajectories that share early prefixes but diverge afterward. The branching is achieved by injecting stochastic perturbations into the SDE transition, with a hyperparameter $s$ controlling the diversity strength among child nodes. This mechanism balances exploration diversity and stability while keeping the marginal distribution unchanged. The rollout continues until reaching the maximum depth $T$, at which point all leaves are collected for reward evaluation.
Formally, following Xue et al. (2025), the reverse-time dynamics can be written in the SDE form:
@@
$z_{i+1}^{(b)}=\mu_{\theta}(z_{i},t_{i})\;+\;g_{t_{i}}\sqrt{h_{i}}\,\xi_{b},\qquad\xi_{b}=\frac{\xi_{0}+s\,\eta_{b}}{\sqrt{1+s^{2}}},\quad b=1,\dots,K,$ (2)
@@
Figure 3: Left: branch rollout process . Middle: leaf rewards are fused upward. Right: depth-wise normalization and pruning yield dense advantages and reduce cost.
@@ 3.3 Reward Fusion and Depth-wise Advantage Estimation
Branch rollouts form a trajectory tree with shared ancestral prefixes, allowing internal node values to be expressed by descendant rewards and enabling backward propagation of leaf signals. However, existing GRPO variants use a single terminal reward at every step, ignoring intermediate states and yielding high-variance, unreliable credit assignment (Fig. 3). BranchGRPO addresses this by propagating leaf rewards upward and converting them into dense step-level advantages via path-probability fusion and depth-wise normalization.
@@
We design a dynamically adjustable reward fusion scheme that aggregates leaf rewards into internal node values through a soft weighting mechanism. For an internal node $n$ with descendant leaves $\mathcal{L}(n)$,
@@
Here $p_{\text{beh}}$ is the behavior policy and $\beta$ controls concentration. $\beta=0$ reduces to uniform averaging; when $\beta=1$, the fusion reduces to weighting by the behavior policy distribution $p_{\text{beh}}(\ell\mid n)$. Uniform averaging is robust to log-prob calibration errors and encourages exploration by retaining low-probability leaves, but introduces variance when branches are many. Path-weighting reduces variance and stabilizes updates, though it may over-concentrate on high-likelihood leaves in deep trees. We empirically compare both variants in Sec. 4.4.
@@
We optimize the standard clipped GRPO loss over tree edges:
where an edge $e$ denotes a transition $(s_{t},a_{t})$ at depth $t$, $\mathcal{E}$ is the set of such edges in a behavior tree, and $\rho_{e}(\theta)=\pi_{\theta}(a_{t}\mid s_{t})/\pi_{\text{old}}(a_{t}\mid s_{t})$.
@@ 3.4 Pruning Strategies
Width Pruning. After computing rewards and advantages for all leaves $\mathcal{L}$, we restrict gradient updates to a subset of them. We investigate two modes. The first, Parent-Top1, retains the child with the higher reward from each parent at the last branching step. This strategy roughly halves gradient computation while ensuring coverage of all parents, yielding stable but slightly less diverse updates. The second, Extreme selection, preserves both the globally best and worst $b$ leaves by reward score. This explicitly maintains strong positive and negative signals, which may enhance exploration but also increase variance.
Depth Pruning. Branch rollouts generate dense rewards across all denoising steps, but computing gradients at every depth remains costly. To improve efficiency, we introduce depth pruning, which skips gradient computation at selected timesteps . Concretely, we maintain a set of pruned depths $\mathcal{D}$ and ignore gradients from nodes at these layers. To prevent permanently discarding certain steps, we adopt a sliding window schedule: the pruned depths gradually shift toward later timesteps as training progresses, until reaching a predefined maximum depth. Formally, pruning is active throughout training, and at fixed intervals the window slides one step deeper until the stop depth is reached.
@@
17:   Depth-wise normalization: standardize per depth, assign edge advantages $A(e)$
18:   Pruning: select nodes for backprop only
19:   Form edge set $\mathcal{E}(c)$ from tree $\mathcal{T}$
20:   Compute $J(\theta)$ (clipped-GRPO over $e\!\in\!\mathcal{E}(c)$, averaged by $|\mathcal{E}(c)|$)
21:   Update policy: $\theta\leftarrow\theta+\eta\,\nabla_{\theta}J(\theta)$
@@ 4.1 Experiment Setup
We evaluate BranchGRPO on HPDv2.1 (Wu et al., 2023) (103k training and 400 balanced test prompts). The backbone is FLUX.1-Dev (Black Forest Labs, 2024), baselines include DanceGRPO and MixGRPO under identical settings. We report efficiency (NFE, iteration time) and quality (HPS-v2.1, PickScore (Kirstain et al., 2023), ImageReward (Xu et al., 2023)), Unified Reward (Wang et al., 2025).
@@ 4.2 Implementation Details
We set the tree depth to $d=20$ and the branch factor to $K=2$, yielding $16$ leaves per rollout before pruning. The branching steps $\mathcal{B}$ use three presets: Dense $(0,3,6,9)$ as the default, Mixed $(0,4,8,12)$, and Sparse $(0,5,10,15)$. The branch correlation sweeps $s\in\{1,2,4,8\}$. Training runs for 300 optimizer steps with gradient accumulation $g=12$ and per-GPU batch size =2=2, on $16\times$ NVIDIA H200 GPUs. Optimization uses AdamW (learning rate $1\times 10^{-5}$, weight decay $1\times 10^{-4}$) with bf16 precision and EMA weights stored on CPU. All GRPO-related hyperparameters are kept identical across methods, with full details deferred to the supplementary material.
@@
FLUX - - - 0.313 0.227 1.112 3.370
DanceGRPO(tf=1.0) 20 20 698 0.360 0.234 1.612 3.388
DanceGRPO(tf=0.6) 20 12 469 0.353 0.228 1.517 3.362
MixGRPO (20,5) 20 5 289 0.359 0.228 1.594 3.380
BranchGRPO 13.68 13.68 493 0.363 0.229 1.603 3.386
BranchGRPO-WidPru 13.68 8.625 314 0.364 0.231 1.609 3.383
BranchGRPO-DepPru 13.68 8.625 314 0.369 0.235 1.625 3.404
BranchGRPO-Mix 13.68 4.25 148 0.363 0.230 1.598 3.384
@@ 4.3 Main Results
Table 1 summarizes efficiency and alignment performance. BranchGRPO consistently outperforms baselines across human-preference metrics while offering favorable compute trade-offs. In particular, BranchGRPO-DepthPruning achieves the best overall alignment, raising HPS-v2.1 from 0.360 (DanceGRPO) to 0.369 and delivering the highest PickScore (0.231), ImageReward (1.625), and Unified Reward (3.404). BranchGRPO-WidthPruning and BranchGRPO-Mix further reduce iteration time to 314s and 148s, respectively, with only marginal drops in quality—making them highly practical for large-scale training. Compared with MixGRPO (289s, HPS-v2.1=0.359, Unified Reward=3.380), BranchGRPO variants yield both stronger alignment and more flexible scaling.
Reward curves in Figure 1 confirm these findings: DanceGRPO(tf=0.6) suffers from instability, while the full-timestep variant converges more smoothly but at high cost. BranchGRPO achieves faster early reward growth, smoother convergence, and higher final rewards. Qualitative comparisons in Figure 4 further show that our method produces sharper details and better semantic alignment than Flux and DanceGRPO.
@@ 4.4 Ablation Studies
Figure 5: Ablation studies of BranchGRPO. Moderate branch correlation, early and denser splits improve reward growth; path-weighted fusion enhances stability; depth pruning achieves the best final reward; and the hybrid ODE–SDE provides the fastest training speed while remaining stable.
Reward Fusion Strategies. Figure 5(d) compares uniform averaging ($\beta=0$) with path-probability weighting ($\beta=1$) under identical training settings. Uniform averaging shows higher variance and a clear late-stage plateau, whereas path-weighted fusion delivers consistently higher and more stable rewards throughout training. This empirically supports Sec. 3.3: uniform averaging is exploration-friendly but noisy, while path weighting aligns credit assignment with the behavior distribution and improves convergence at no extra cost.
Pruning Strategies. Figure 5(e) compares pruning methods applied after depth-wise normalization and only during backpropagation. For depth pruning, we adopt a sliding-window schedule over denoising steps, the window is initialized at the last split point, has a fixed size of $4$, and shifts by one denoising step every $30$ training iterations. GRPO losses and gradients in the active window are skipped, while forward sampling remains unchanged. This schedule yields the best final reward and reveals substantial redundancy at late timesteps. For width pruning (Parent-Top1), we retain only the locally better child at each branch for gradient updates, effectively halving updates and producing the smoothest, lowest-variance curve, though with slightly lower final reward than depth pruning. Width pruning (Extreme-$b$) keeps both the globally best and worst $b$ leaves, injecting stronger positive/negative signals and remaining competitive at the end, but with higher variance.